OpenStudy (anonymous):

What is the solution to the equation? 2q + 18 = –5q – 3 A. q = –5 B. q = –3 C. q = 3 D. q = 5

OpenStudy (anonymous):

medal if helped

OpenStudy (anonymous):

pls

OpenStudy (anonymous):

heelp

OpenStudy (elonasushchik):

2q + 18 = -5q - 3 +5q +5q 7a + 18 = -3 -18 -18 7a = -21 --------- 7 7 a = -3

OpenStudy (elonasushchik):

Medal?! Thx

OpenStudy (anonymous):

thanks sooooooooooooooooooo much

OpenStudy (elonasushchik):

Your welcome:))

OpenStudy (anonymous):

oh but its q not a

OpenStudy (anonymous):

is it q=-3

OpenStudy (elonasushchik):

It's -3 though...

OpenStudy (anonymous):

oh ok thanks

OpenStudy (elonasushchik):

Np!